OpenStudy (anonymous):

Pls help. A box contains 50 rose cuttings out of that 20 red, 15 yellow, 5 white and 10 purple. If 6 rose cuttings are selected randomly a.P(5 red rose cuttings) b.P(2 red, 2 yellow, 2 purple) c.(<3 purple)

OpenStudy (anonymous):

\[P(x) = ^nC_x p^x (1 - p)^{n - x}\] where p is the success rate, n is number of trials and x is the amount wanted a.P(5 red rose cuttings) p=20/50=2/5 n=6 x=5 \[P(5 red) = ^6C_5 0.4^5 (1 - 0.4)^{6 - 5}\]\[= 6* 0.4^5 *0.6^1\]\[=0.036864\]

OpenStudy (anonymous):

We use the same method to work out the separate probabilities for b and c. b. P(2 red, 2 yellow, 2 purple)=P(2red)*P(2yellow)*P(2purple) c. P(<3 purple)=P(0purple)+P(1purple)+P(2purple)

OpenStudy (anonymous):

For the other two I get b. P=0.00992 and c. P=0.98415

OpenStudy (anonymous):

Is this right